LPC meeting summary 25-04-2016 - final

Minutes overview      LPC home

Minutes and Summary

Main purpose of the meeting: Commissioning and first Stable Beams.

Introduction (Jamie Boyd)

Jamie shortly summarised the status of the machine and gave an overview on the first four fills with Stable Beams. These fills were slightly anticipated to the end of the previous weeks due to the fast progress of the LHC commissioning. This week will be devoted to scrubbing. As it looks the de-conditioning of the machine is not too bad.

Jamie re-iterated on the filling schemes for the first intensity ramp-up steps. All schemes are available via the LPC web page. ATLAS would prefer short trains of 12 bunches for the O(80b) step. The LPC will look into this and see if a scheme can be generated which does not penalize too much LHCb and ALICE.
It was announced that via the LPC web page a service will be provided, which allows to download the csv file for each fill which reached Stable Beams. The service will be URL based and is intended to be used by automatic scripts.

It was asked if experiments are ready to produce Massi files for the physics fills. Witold remarked that the current Massi files of CMS might need and update.

The discussion on the insertion of the CTPPS pots is summarised in the CTPPS section below.

ATLAS (Alessandro Cerri)

Alessandro summarised briefly the commissioning activities of ATLAS during the first Stable Beams. ATLAS observed some drift of the beamspot in the z-coordinate which was discussed briefly: It was pointed out that the OFB (Orbit Feedback system) was not switched on at during the first part of one of the fills and that once it was switched on by the shift crew, the z-position of the beam spot jumped back to its initial value. Jörg Wenninger pointed out that the crossing angle in ATLAS would vary if the OFB is not switched on and this would translate into a drifting in z of the beam spot. Jörg also pointed out that in a long fill CMS which has a horizontal crossing plane, would always have a slower luminosity decay than ATLAS due to the fact that the vertical beam-size shrinks during a long fill.

ATLAS reminded to schedule the loss maps for the VdM scan not directly before the scan in order not to loose the radiation dependent memory of the BCM detectors.

ATLAS asked about the restrictions for µ-separation with the new 40cm optics. Jörg replied that this was not yet discussed among experts. A total separation of 3σ should be acceptable, but if experiments want to go to a µ of 0.1 more than 4σ would be required. This issue should be followed up by the LPC with the machine experts.

AFP (Maciej Marek Trzebinski)

Maciej summarised the insertion and ongoing qualification of the AFP pots. The insertions during the ramp up will be done as outlined here:


AFP would be interested in a low pileup (µ≈0.1) fill during their commissioning activities. This would have to be achieved with beam separation and as mentioned above this needs to be discussed among experts.

During one of the insertions the AFP pot spontaneously went back to the "Home" position. It is not yet understood what happened and unfortunately the relevant data to investigate post mortem is not logged by AFP. It was speculated that the LVDT readings could have gone above warning level which is worrisome since the dump level is only 100µm above the warning level. Sune Jakobsen mentioned that in ALFA such problems have been seen and the readout algorithms have been improved to address these fluctuations in the LVDT readings.

ALFA (Sune Jakobsen)

Sune was presenting the radiation related observations made in ALFA during the AFP insertion. ALFA was seeing variations of their trigger rate when TCLs or the AFP pots were inserted. However the sensitivity of the newly installed BLMs is not high enough to give clear signals with the radiation levels during the fills with 3 and 12 bunches. It was concluded that more data with higher intensity beams is needed to make some quantitative statements on the radiation levels at ALFA due to the presence of AFP.

CMS (Silvia Goy Lopez)

Silvia summarised the commissioning activities of CMS with first collisions and with first Stable Beams. She pointed out that during the Quiet Beams period, CMS observed a background spike which was not crossing any threshold yet, but was at a level which would have prevented CMS from switching on the Tracker, if this background level would have been constantly present at the start of the Quiet Beam period. CMS is asking for a clear definition of the Quiet Beam conditions. It was pointed out by Jörg that Quiet Beams means that the conditions are the same as for Stable Beams but the the validation of the settings has not yet been completed (e.g. no complete set of loss maps yet)

CMS pointed out that any kind of collisions are very interesting for the experiments commissioning and therefore it would be nice to have a clear message on page one when these collisions are happening, so that also shift crews with limited expertise understand that they are happening. In the same spirit it would be nice to clearly announce on page 1 when losses are expected.

Silvia pointed out that CMS will need to go through some more commissioning steps once the magnetic field will be at 3.8T.

A list of requests of CMS was presented:


LHCb (Patrick Robbe)

All detectors are now aligned in time and space and LHCb is ready for data taking in 2016




CTPPS (Mario Deile)

Mario summarised the operations of CTPPS and Totem Roman Pots which includes the alignment of all pots, loss maps with the CTPPS pots at 15σ without margin, and the first insertion in a fill with Stable Beams. On the 23rd of April the pots were inserted to 15σ without margin during Stable Beams with 3 circulating bunches. On the 25th of April they were inserted to 15σ with 0.5mm margin into a fill with 12 bunches (during Stable Beams). Mario shortly discussed the first data taken during these insertions.

Mario presented the agreed plan (with MPP, OP and the collimator team) of the CTPPS insertion during the ramp up phase:


ALICE (Siegfried Foertsch)

Siegfried summarised the commissioning activities of ALICE during the first fills with Stable Beams. The beams in ALICE are still separated "by hand" but the automated procedure is being developed.

Siegfried pointed out that ALICE would run with an external crossing angle of 160µrad and +/+ polarity during all VdM scans.


Background tests with increased gas pressure (Reyes Alemany Fernandez)

The background tests with increased gas pressure around IP1 and IP5 (organised by Reyes) should be done in a fill with low number of bunches (50-80 bunches). This test should not be done at the same time as the afterglow studies requested by ATLAS. The LPC should follow up on the organisation of this test.

Mini-scan for a lumi pre-calibration (CMS/ATLAS inter-calibration) (Witold Kozanecki)

Witold asked if there was an interest to inter-calibrate the luminosity of the experiments (especially those of ATLAS and CMS) with an early Mini-scan. This would take of the order of 45min per experiment and would be done during a physics fill. The analysis would be expected to take of the order of three days.

In a short discussion it became clear that CMS would be moderately interested in such an exercise but the resulting calibration would not be used for the online lumi publication in subsequent fills

It was concluded that experiments should come back with a clear statement on this in the next LPC meeting.